A salient pole synchronous generator has the following per unit parameters;xd=1.2,xq=0.8 ra=00.025. compute the excitation voltage Ef on a per unit basis,when the generator is delivering rated kva at rated voltage and at power factor of 0.8 lagging and 0.8 leading


1
Expert's answer
2021-02-09T07:08:03-0500

"Xd= 1.2 pu \newline
Xq= 0.8 pu\newline
Ra= 0.025 pu\newline"


As it's given that rated kva and rated voltage. We have

"Terminal\space voltage \space V=1pu\newline
Armature \space current \space=\space1pu"


1) Power factor "= 0.8 lag"

This implies Power factor angle ", \phi=cos^{-1}(0.8) = 36.87\degree"


We have,


"Tan\psi= \frac{Vsin\phi+IaXq}{Vcos\phi+IaRa}=\frac{0.6+0.8}{0.8+0.025}=1.697."

This implies "\psi=Tan^{-1}(1.697) =59.49\degree."


Load angle "(\delta) =\psi-\phi=59.49-36.87=22.62\degree"


We have,

"Id =Iasin\psi=sin59.49=0.86\newline Iq=Iacos\phi= cos59. 49=0.51."


Hence Excitation voltage is,


"E= Vcos\delta+IqRa+IdXd\newline 
E= cos22. 62+0.51\times 0.025+0.86\times1.2\newline
E= 1.97 pu"


There fore "E\angle\delta=1.97\angle22.62 pu"


2) Power factor is 0.8 lead.


This implies Power factor angle ", \phi=cos^{-1}(0.8) = 36.87\degree"


We have,


"Tan\psi= \frac{Vsin\phi-IaXq}{Vcos\phi+IaRa}=\frac{0.6-0.8}{0.8+0.025}=-0.242"

This implies "\psi=Tan^{-1}(-0.242) =-13.63\degree."


Load angle "(\delta) =\psi+\phi=-13.63+36.87=23.24\degree"


We have,

"Id =Iasin\psi=sin13.63=0.24\newline Iq=Iacos\phi= cos13.63=0.97"


Hence Excitation voltage is,


"E= Vcos\delta+IqRa-IdXd\newline 
E= cos23.24+0.97\times 0.025-0.24\times1.2\newline
E= 0.66pu"


There fore "E\angle\delta=0.66\angle23.24pu"